# https://sinhasaptarshi.github.io/escounts

### v1

@@ Abstract
Video repetition counting infers the number of repetitions of recurring actions or motion within a video. We propose an exemplar-based approach that discovers visual correspondence of video exemplars across repetitions within target videos. Our proposed Every Shot Counts (ESCounts) model is an attention-based encoder-decoder that encodes videos of varying lengths alongside exemplars from the same and different videos. In training, ESCounts regresses locations of high correspondence to the exemplars within the video. In tandem, our method learns a latent that encodes representations of general repetitive motions, which we use for exemplar-free, zero-shot inference. Extensive experiments over commonly used datasets (RepCount, Countix, and UCFRep) showcase ESCounts obtaining state-of-the-art performance across all three datasets. On RepCount, ESCounts increases the off-by-one from $0.39$ to $0.56$ and decreases the mean absolute error from $0.38$ to $0.21$. Detailed ablations further demonstrate the effectiveness of our method.
@@ 1 Introduction
Recent years have come with tremendous progress in video understanding. Large Visual Language Models (VLMs) have been adopted for many vision tasks including video summarisation [44, 32, 60], localisation [45, 57], and question answering (VQA) [1, 21, 39, 64]. Despite their great success, recent analysis [24] shows that VLMs can still fail to correctly count objects or actions. Robust counting can be challenging due to appearance diversity, limited training data, and most critically semantic ambiguity – identifying ‘what’ exactly to count.
Evidence in developmental psychology and cognitive neuroscience [50, 53, 54] shows that infants fail to differentiate the number of hidden objects when they are not shown and counted to them first, suggesting an upper limit of individual objects that can be represented in working memory. However, infants that were first exposed to an instance of the object could better approximate the carnality. This shows that counting is a visual exercise of matching to given exemplars, and is developed in humans prior to understanding their semantics.
Object-counting in images has recently exploited exemplars to improve performance [38, 10]. In training, models attend to one or more exemplars of ‘what’ object(s) to count alongside learnt embeddings for exemplar-free counting. During inference, only the learnt embeddings are used, in a zero-shot manner, without knowledge of exemplars. As videos are of variable length and repetitions vary in duration, these approaches are not directly applicable to counting in videos.
@@
Taking inspiration from image-based approaches, we address Video Repetition Counting (VRC) using exemplars, for the first time. We differ from prior works that formulate VRC as classifying a preset number of repetitions [6, 16, 67], or consecutively detecting relevant parts of the repetition [15, 20, 30], such as the start/end. Instead, we argue that learning correspondences to reference exemplar(s) can provide a strong prior for discovering repetitions in videos. We propose Every-Shot Counts (ESCounts), a transformer-based encoder-decoder that encodes videos of varying lengths alongside exemplars and learns representations of general repeating motions, as shown in Figure 1. Similar to prior works [20, 30], we use the density map to regress the temporal location of each repetition. During training, we learn an exemplar latent representation, which we use for inference where no exemplars are available.
In summary, our contributions are as follows: (i) We introduce exemplar-based counting for VRC (ii) We propose an attention-based encoder-decoder that learns to correspond exemplars to a query video of varying length, and predict repetitions. (iii) We use a learnt latent for general motions in repetitions enabling our approach to predict the number of repetitions during inference without exemplars, (iv) We evaluate our approach on the three datasets commonly used for repetition counting in videos: RepNet [20], Countix [16], and UCFRep [65]. Our approach achieves a new state-of-the-art in every benchmark, even on Countix where the start-end times of repetitions are not annotated.
@@ 2.1 Object Counting in Images
Class-Specific Counting. These methods learn to count objects of singular classes or sets of categories e.g. people [56, 29], cars [19], wildlife [4]. A large portion of object-counting approaches [19, 56, 12, 40, 47] have relied on detecting target objects and counting their instances. Traditional methods have used hand-crafted feature descriptors to detect human heads [56] or head-shoulders [29] for crowd-counting. Other methods have used blobs [26], individual points [34], and object masks [13] for detecting and counting instances. Though object detection can be a preliminary step before counting, detection methods rely strongly on the performance of the object detector which can be less effective in densely crowded images [12]. A different set of methods has instead relied on regression, to either regress to the target count [12, 61] or estimate a density map [27, 41, 66].
Class-Agnostic Counting. Class-specific counting approaches are impractical for general setting, as prior knowledge of the object category is not available. Recent works [10, 3, 31, 48] have used one (or a few) exemplars as references to estimate a density map for unknown target classes. Exemplars for object counting were first introduced by Lu et al. [38]. Building on the property of image self-similarity[38] proposed a convolutional matching network and cast counting as an image-matching problem, where exemplar patches from the same image are used to match against other patches within the image. Following up, Liu et al. [10] proposed a transformer-based feature interaction module. They used an encoder for the query image and a convolution-based encoder for the exemplar. The interaction module cross-attended information between the exemplar and image and used a convolutional decoder to regress the density map. Recent approaches have also fused text and visual embeddings [3], used contrastive learning across modalities [24], and generated exemplar prototypes using stable diffusion [62]. Inspired by these methods, we propose a fully-attention-based encoder-decoder that extends exemplar-based counting to VRC. Our approach is invariant to video lengths and can use both learnt or encoded exemplars.
@@ 2.2 Video Repetition Counting (VRC)
Compared to image-based counting, video repetition counting has been less explored. Early approaches have compressed motion into a one-dimensional signal and recovered the repetition structure from the signal’s period [2, 25, 37, 42]. The periodicity can then be counted by Fourier analysis [5, 7, 14, 43, 2, 25], peak detection [52], or wavelet analysis [46]. However, these methods are limited to uniformly periodic repetitions. For non-periodic repetitions, localisation frameworks [11, 36, 49, 33] have been adapted. Zhang et al. [65] proposed a context-aware scale-insensitive framework to count repetitions of varying scales and duration. Their method exhaustively searched for pairs of consecutive repetitions followed by a prediction refinement module. Recent methods [15, 16, 20] have also extended image self-similarity to the temporal dimension with temporal self-similarity matrices (TSM). TSM is constructed using pair-wise similarity of embeddings over temporal locations. RepNet [16] used a transformer-based period predictor. To count repetitions with varying speeds, Trans-RAC [20] modified TSM to use multi-scale sequence embeddings. For counting under poor lighting conditions, [67] used audio in tandem with video within a multi-modal framework. They selectively aggregated information from the two modalities using a reliability estimation module. Li et al. [30] also used multi-modal inputs with optical flow as an additional signal supporting RGB for detecting periodicity. Yao et al. [63] proposed a lightweight pose-based transformer model that used action-specific salient poses as anchors and matched them across the video. However, the requirement of salient pose labels for each action limits its generalisability to unseen repetitions.
The above methods do not utilise the correspondences discovered by exemplar repetitions. Thus, do not relate variations in the action’s performance, alongside other visual variations such as viewpoint and camera motion. To handle such cases, we propose using action exemplars as references for VRC. Exemplars have previously been used in videos for action recognition tasks [17, 23, 59, 58]. [58] used silhouette/pose exemplars for classifying action sequences into predefined action categories using a Bayes classifier. [59] converted training videos to visual vocabulary and used the most discriminative visual words as exemplars for detecting actions. These methods are limited to a predefined set of classes and will be ineffective for class-agnostic tasks such as repetition counting. To our knowledge, we are the first to use exemplars for repetition counting in videos.
@@ 3 Every Shot Counts (ESCounts) Model
In this section, we introduce our ESCounts model (overviewed in Figure 2). We formally define encoding variable length videos alongside our model’s output, in Section 3.1. We then introduce the attention-based decoder that corresponds the input video to exemplars in Section 3.2. Predictions over temporally shifted inputs are combined for a final prediction, detailed in Section 3.3.
@@ 3.1 Input Encoding and Output Prediction
We denote the full video as $\mathbf{v}$ of varying $\mathcal{T}$ length and a fixed $H\times W$ spatial resolution. Segment $\mathbf{e}_{s}$ containing a single instance of the repeating action we wish to count, is selected as an exemplar. Exemplars are defined based on provided [[[[start, end]]]] labels of every repetition in the video [^1] For datasets where the start/end times are not available, pseudo-labels are used instead by uniformly dividing the video given the ground truth count.. During training, we select one or more exemplar shots $\mathcal{S}\subseteq\{\mathbf{e}_{1},\dots,\mathbf{e}_{s}\}$. Each training sample is thus a combination of the query video and a set of exemplars $(\mathbf{v},\mathcal{S})$.
We tokenise and encode the video $\mathbf{v}$ from its original size $\mathcal{T}\times H\times W$ into spatiotemporal latents $\mathbf{z}_{v}$. Encoder $\mathcal{E}$ is applied over a fixed-size sliding window, to account for the video’s variable length. The encoded video is represented by $\mathbf{z}_{v}\in\mathbb{R}^{M\times C}$ of $M=\mathcal{T^{\prime}}H^{\prime}W^{\prime}$ spatiotemporal resolution with $C$ channels. We note that $M$ is not a fixed number, as it depends on the video’s length $\mathcal{T}$. We add sinusoidal positional encoding to account for the relative order of these spatiotemporal latents while accommodating the variable video length.
During training, we select exemplars $\mathcal{S}$ from either the same video or another video containing the same action label; e.g. given a video containing push-up actions, we can sample exemplars from other videos showcasing the same action within the training set. We define a probability $p$ of sampling the exemplar from a different video; i.e. $p=0$ implies exemplars are only sampled from the same video, whereas for $p=1$ exemplars are always sampled from another video [^2] We ablate $p$ in our experiments.. In training, exemplars are sampled randomly from the labelled repetitions of the video. We use $\mathcal{E}$ to encode latent representations from each exemplar $\mathbf{e}_{s}\in\mathcal{S}$. We use the same encoder $\mathcal{E}$ for encoding $\mathbf{v}$ and $\mathbf{e}_{s}$ to allow correspondence.
We construct the ground truth density map $\mathbf{d}$ from the labelled repetitions in the video as a 1-dimensional vector. To match the downsampled temporal resolution of our input video $\mathcal{T^{\prime}}$, we also temporally downsample the ground-truth labels. The density map takes low values ($\approx 0$) at temporal locations without repetitions in the video, and high within repetitions. We use a normal distribution $\mathcal{N}$ centred around each repetition with $(\mu_{i}=\frac{t_{s}+t_{e}}{2},\sigma)$, where $t_{e}$ and $t_{s}$ are the start and end times of each repetition $i$.
@@ 3.2 Latent Exemplar Correspondence
Given both the encoded video $\mathbf{z}_{v}=\mathcal{E}(\mathbf{v})$ and exemplars $\mathbf{z}_{s}=\mathcal{E}(\mathbf{e}_{s})\;\forall\;\mathbf{e}_{s}\in\mathcal{S}$ latents, we use an attention-based decoder $\mathcal{D}(\mathbf{z}_{v},\mathbf{z}_{s})$ to attend to every repetition in the video $v$ that matches the encoded exemplar. Decoder $\mathcal{D}$ takes the encoded video $\mathbf{z}_{v}$ as input and predicts the location of every repetition in the video. The decoder outputs a 1-dimensional predicted density map of length $\mathcal{T}^{\prime}$ corresponding to the occurrences of the repeating action given the exemplars.
@@
Cross-attention Blocks. We explore the similarity between representations of exemplars and the query video to predict the corresponding locations of repetitions that match the exemplar. Thus, inspired by [10], we use cross-attention to relate exemplar and video latents. We define $L$ cross-attention blocks. Each block initially self-attends $\text{SA}(\cdot)$ the video latents $\mathbf{z}_{l}\in\mathbb{R}^{M\times C}$ with multi-head self-attention. We note that for the first layer, $\mathbf{z}_{1}=\mathbf{z}_{v}$. Then we relate exemplar and video information by cross-attending $\text{CA}(\cdot)$ video and exemplar latents. The block’s initial self-attention operation is formulated as:
@@
where $\text{LN}(\cdot)$ is layer normalisation.
@@
We also want to learn representations for repeating motions to estimate repetitions without explicitly providing exemplars. We thus define a learnable latent $\mathbf{z}_{0}$ to cross-attend $\mathbf{z}_{v}$. At each training iteration, we select exemplars from $\{\mathbf{z}_{0},\mathbf{z}_{1},\text{...},\mathbf{z}_{s}\}$ and perform CA with $\mathbf{z}_{0}$ or $\{\mathbf{z}_{1},\text{...},\mathbf{z}_{s}\}$. At inference, we use $\mathbf{z}_{0}$.
We use a multi-layer perceptron MLP on the exemplar-fused latents $\mathbf{z}^{\prime\prime}_{l}$.
@@
The cross-attention blocks’ output is defined as $\mathbf{z}_{L}\in\mathbb{R}^{M\times C}$.
Window Self-attention Blocks. We explore the spatio-temporal inductive bias within the self-attention blocks. Each latent thus attends locally, to its spatio-temporal neighbouring tokens, over $L^{\prime}$ window self-attention $\text{WSA}(\cdot)$ [35] layers. We denote $\forall\,l\in\,\{L+1,\dots,L+L^{\prime}\}$ :
@@
The output of the window self-attention blocks is of size $\mathbf{z}_{L+L^{\prime}}\in\mathbb{R}^{M\times C}$ In turn, $\mathbf{z}_{L+L^{\prime}}$ encodes repetition-relevant features over space and time. Thus, we use $\mathbf{z}_{L+L^{\prime}}$ to predict density map $\tilde{\mathbf{d}}$ for the occurrences of the target repeating action over time. We use a fully connected layer to project the latent to a 1-channel vector, i.e. MLP: ${\mathbb{R}^{M\times C}\rightarrow\mathbb{R}^{M}}$. We then pool the spatial resolution $H^{\prime}W^{\prime}$ resulting to predicted density map $\tilde{\mathbf{d}}\in\mathbb{R}^{\mathcal{T}^{\prime}}$.

### v2

@@ Abstract
Video repetition counting infers the number of repetitions of recurring actions or motion within a video. We propose an exemplar-based approach that discovers visual correspondence of video exemplars across repetitions within target videos. Our proposed Every Shot Counts (ESCounts) model is an attention-based encoder-decoder that encodes videos of varying lengths alongside exemplars from the same and different videos. In training, ESCounts regresses locations of high correspondence to the exemplars within the video. In tandem, our method learns a latent that encodes representations of general repetitive motions, which we use for exemplar-free, zero-shot inference. Extensive experiments over commonly used datasets (RepCount, Countix, and UCFRep) showcase ESCounts obtaining state-of-the-art performance across all three datasets. Detailed ablations further demonstrate the effectiveness of our method.
@@ 1 Introduction
In recent years, tremendous progress has been made in video understanding. Visual Language Models (VLMs) have been adopted for many vision tasks including video summarisation [49, 36, 65], localisation [50, 62], and question answering (VQA) [1, 23, 44, 69]. Despite their great success, recent analysis [27] shows that VLMs can still fail to count objects or actions correctly. Robust counting can be challenging due to appearance diversity, limited training data, and the semantic ambiguity of identifying ‘what’ to count.
Evidence in developmental psychology and cognitive neuroscience [56, 59, 60] shows that infants fail to differentiate the number of hidden objects if not shown and counted to them first, suggesting an upper limit of individual objects in working memory. However, infants exposed to an instance of the object first could better approximate carnality. This shows that counting is a visual exercise of matching to exemplars, and is developed before understanding their semantics.
Object-counting in images has recently exploited exemplars to improve performance [43, 38]. In training, models attend to one or more exemplars of ‘what’ object(s) to count alongside learnt embeddings for exemplar-free counting. During inference, only the learnt embeddings are used for zero-shot counting without knowledge from exemplars. As videos are of variable length and repetition durations vary, these approaches are not directly applicable to videos.
@@
Taking inspiration from image-based approaches, we address Video Repetition Counting (VRC) with exemplars for the first time. We differ from prior works that formulate VRC as classifying a preset number of repetitions [6, 15, 72], or detecting relevant parts (start/end) of repetitions [14, 22, 33]. Instead, we argue that learning correspondences to reference exemplar(s) during training can provide a strong prior for discovering correspondences across repetitions at inference. We propose Every-Shot Counts (ESCounts), a transformer-based encoder-decoder that during training encodes videos of varying lengths alongside exemplars and learns latents of general repeating motions, as shown in Fig. 1. Similar to [22, 33], we use density maps to regress the temporal location of repetitions. At inference, learnt latents are used for exemplar-free counting.
In summary, our contributions are as follows: (i) We introduce exemplar-based counting for VRC (ii) We propose an attention-based encoder-decoder that corresponds exemplars to a query video of varying length. (iii) We learn latents for general repeating features and use them to predict the number of repetitions during inference without exemplars, (iv) We evaluate our approach on the three commonly-used VRC datasets: RepNet [22], Countix [15], and UCFRep [70]. Our approach achieves a new state-of-the-art in every benchmark, even on Countix where start-end times of repetitions are not annotated.
@@ 2.1 Object Counting in Images
Class-specific counting. These methods learn to count objects of singular classes or sets of categories e.g. people [61, 32], cars [21], or wildlife [4]. A large portion of object-counting approaches [21, 61, 11, 45, 52] have relied on detecting target objects and counting their instances. Traditional methods have used hand-crafted feature descriptors to detect human heads [61] or head-shoulders [32] for crowd-counting. Other methods have used blobs [29], individual points [39], and object masks [12] for detecting and counting instances. Though object detection can be a preliminary step before counting, detection methods rely strongly on the object detector’s performance which can be less effective in densely crowded images [11]. Other methods instead relied on regression, to either regress to the target count [11, 66] or estimate a density map [30, 46, 71].
Class-agnostic counting. Class-specific counting approaches are impractical for general settings where prior knowledge of the object category is not available. Recent works [38, 3, 35, 54] have used one (or a few) exemplars as references to estimate a density map for unknown target classes. Building on the property of image self-similarity[43] proposed a convolutional matching network. They cast counting as an image-matching problem, where exemplar patches from the same image are used to match against other patches within the image. Following up, Liu et al. [38] used an encoder for the query image, a convolution-based encoder for the exemplar, and an interaction module to cross-attend information between the exemplar and the image. A convolutional decoder was used to regress the density map. Recent approaches have also fused text and visual embeddings [3], used contrastive learning across modalities [27], and generated exemplar prototypes using stable diffusion [67]. Inspired by these methods, we propose an attention-based encoder-decoder that extends exemplar-based counting to VRC. Our approach is invariant to video lengths and can use both learnt or encoded exemplars.
@@ 2.2 Video Repetition Counting (VRC)
Compared to image-based counting, video repetition counting has been less explored. Early approaches have compressed motion into a one-dimensional signal and recovered the repetition structure from the signal’s period [2, 28, 42, 47]. The periodicity can then be counted with Fourier analysis [5, 7, 13, 48, 2, 28], peak detection [58], or wavelet analysis [51]. However, these methods are limited to uniformly periodic repetitions. For non-periodic repetitions, temporal understanding frameworks [10, 41, 55, 37] have been adapted. Zhang et al. [70] proposed a context-aware scale-insensitive framework to count repetitions of varying scales and duration. Their method exhaustively searches for pairs of consecutive repetitions followed by a prediction refinement module. Recent methods [14, 15, 22] have also extended image self-similarity to the temporal dimension with Temporal Self-similarity Matrices (TSM). TSM is constructed using pair-wise similarity of embeddings over temporal locations. RepNet [15] used a transformer-based period predictor. To count repetitions with varying speeds, Trans-RAC [22] modified TSM to use multi-scale sequence embeddings. For counting under poor lighting conditions, [72] used both audio and video in a multi-modal framework. They selectively aggregated information from the two modalities using a reliability estimation module. Li et al. [33] also used multi-modal inputs with optical flow as an additional signal supporting RGB for detecting periodicity.
Recent works attempt to utilise spatial [68] or temporal [73, 34] saliency for repetition counting. Yao et al. [68] proposed a lightweight pose-based transformer model that used action-specific salient poses as anchors. The need for salient pose labels for each action limits generalisability to unseen repetitions. Zhao et al. [73] used a dual-branch architecture to first select repetition-relevant video segments and then attend over these frames. Li et al. [34] used a joint objective to localise and binary classify regions as (non-)repetitive.
The above methods do not utilise the correspondences discovered by exemplar repetitions. Thus, do not relate variations in the action’s performance. We propose using action exemplars as references for VRC. Exemplars have previously been used in videos for action recognition tasks [18, 26, 64, 63][63] used silhouette/pose exemplars for classifying action sequences into predefined categories.[64] converted training videos to a visual vocabulary and used the most discriminative visual words as exemplars. These methods are limited to a predefined set of classes. To our knowledge, we are the first to use exemplars for repetition counting in videos.
@@ 3 Every Shot Counts (ESCounts) Model
In this section, we introduce our ESCounts model (overviewed in Fig. 2). We formally define encoding variable length videos alongside our model’s output, in Sec. 3.1. We introduce the attention-based decoder that corresponds the input video to training exemplars and learnt latents in Sec. 3.2. Predictions over temporally shifted inputs are then combined, detailed in Sec. 3.3.
@@ 3.1 Input Encoding and Output Prediction
We denote the full video as $\mathbf{v}$ of varying $\mathcal{T}$ length and fixed $H\times W$ spatial resolution. Segment $\mathbf{e}_{s}$ containing a single instance of the repeating action we wish to count, is selected as an exemplar. Exemplars are defined based on provided [[[[start, end]]]] labels of every repetition in the video [^1] For datasets where the start/end times are not available, pseudo-labels are used instead by uniformly dividing the video by the ground truth count.. During training, we select one or more exemplar shots $\mathcal{S}\subseteq\{\mathbf{e}_{1},\dots,\mathbf{e}_{s}\}$. Each training instance is a combination of the query video and the set of exemplars $(\mathbf{v},\mathcal{S})$.
We tokenise and encode the video $\mathbf{v}$ from its original size $\mathcal{T}\times H\times W$ into spatiotemporal latents $\mathbf{z}_{v}$. To account for the video’s variable length, encoder $\mathcal{E}$ is applied over a fixed-size sliding window. The encoded video is represented by $\mathbf{z}_{v}\in\mathbb{R}^{M\times C}$ of $M=\mathcal{T^{\prime}}H^{\prime}W^{\prime}$ spatiotemporal resolution with $C$ channels. We note that $M$ is not a fixed number, as it depends on the video’s length $\mathcal{T}$. We add sinusoidal positional encoding to account for the relative order of these spatiotemporal latents while accommodating the variable video length.
For training only, we select exemplars $\mathcal{S}$ from either the same video or another video of the same action category; e.g. given a video containing push-up actions, we can sample exemplars from other videos showcasing the same action within the training set. We define a probability $p$ of sampling the exemplar from a different video; i.e. $p=0$ implies exemplars are only sampled from the same video, whereas for $p=1$ exemplars are always sampled from another video [^2] We ablate $p$ in our experiments.. We sample exemplars randomly from the labelled repetitions of the video. We use $\mathcal{E}$ to encode latent representations from each exemplar $\mathbf{e}_{s}\in\mathcal{S}$. We use the same encoder $\mathcal{E}$ for encoding $\mathbf{v}$ and $\mathbf{e}_{s}$ to enable direct correspondence.
We construct the ground truth density map $\mathbf{d}$ from the labelled repetitions in the video as a 1-dimensional vector. To match the downsampled temporal resolution of our input video $\mathcal{T^{\prime}}$, we also temporally downsample the ground-truth labels. The density map takes low values ($\approx\!0$) at temporal locations without repetitions and high values within repetitions. We use a normal distribution $\mathcal{N}$ centred around each repetition with $(\mu_{i}=\frac{t_{s}+t_{e}}{2},\sigma)$, where $t_{e}$ and $t_{s}$ are the start and end times of each repetition $i$.
@@ 3.2 Latent Exemplar Correspondence
Given both the encoded video $\mathbf{z}_{v}=\mathcal{E}(\mathbf{v})$ and exemplars $\mathbf{z}_{s}=\mathcal{E}(\mathbf{e}_{s})\;\forall\;\mathbf{e}_{s}\in\mathcal{S}$, we use an attention-based decoder $\mathcal{D}(\mathbf{z}_{v},\mathbf{z}_{s})$ to learn a correspondence between every repetition in the video $v$ and the encoded exemplar. Decoder $\mathcal{D}$ takes the encoded video $\mathbf{z}_{v}$ as input and predicts the location of every repetition in the video. The decoder outputs a 1-dimensional predicted density map of length $\mathcal{T}^{\prime}$ corresponding to the occurrences of the repeating action given the exemplars.
@@
Cross-attention Blocks. We explore the similarity between exemplars and query video representations to predict the corresponding locations of repetitions that match the exemplar. Thus, inspired by [38], we use cross-attention to relate exemplar and video encodings. We define $L$ cross-attention blocks. Each block initially Self-Attends $\text{SA}(\cdot)$ the video latents $\mathbf{z}_{l}\in\mathbb{R}^{M\times C}$ with multi-head self-attention.
We note that for the first layer, $\mathbf{z}_{1}=\mathbf{z}_{v}$. We then relate exemplar and video by Cross-Attending $\text{CA}(\cdot)$ video and exemplar encodings. The block’s initial self-attention operation is formulated as:
@@
where $\text{LN}(\cdot)$ is Layer Normalisation. It is essential to self-attend across the video first to capture the features of the repeated actions within the video, and enforce feature correspondence between repetitions.
@@
We also want to learn repeating motions to estimate repetitions without explicitly providing exemplars. We thus define a learnable latent $\mathbf{z}_{0}$ to cross-attend $\mathbf{z}_{v}$. At each training step, we select exemplars from $\{\mathbf{z}_{0},\mathbf{z}_{1},\text{...},\mathbf{z}_{s}\}$ and perform CA with $\mathbf{z}_{0}$ or $\{\mathbf{z}_{1},\text{...},\mathbf{z}_{s}\}$. Importantly, at inference, we use only $\mathbf{z}_{0}$.
We obtain the cross-attention blocks’ output, defined as $\mathbf{z}_{l+1}\in\mathbb{R}^{M\times C}$, with a Multi-Layer Perceptron MLP on the exemplar-fused latents $\mathbf{z}^{\prime\prime}_{l}$.
@@
Window Self-attention Blocks. We explore the spatio-temporal inductive bias within the self-attention blocks. For this, each latent attends locally to its spatio-temporal neighbouring tokens, over $L^{\prime}$ Window Self-Attention $\text{WSA}(\cdot)$ [40] layers. We denote $\forall\,l\in\,\{L+1,\dots,L+L^{\prime}\}$ :
@@
The output of the WSA blocks is of size $\mathbf{z}_{L+L^{\prime}}\in\mathbb{R}^{M\times C}$ In turn, $\mathbf{z}_{L+L^{\prime}}$ encodes repetition-relevant features over space and time and is used to predict density map $\tilde{\mathbf{d}}$ for the occurrences of the target repeating action over time. We use a fully connected layer to project the latent to a 1-channel vector, i.e. MLP: ${\mathbb{R}^{M\times C}\rightarrow\mathbb{R}^{M}}$. We then vectorise the spatial resolution $H^{\prime}W^{\prime}$ whilst maintaining $\mathcal{T}^{\prime}$ resulting to the predicted density map $\tilde{\mathbf{d}}\in\mathbb{R}^{\mathcal{T}^{\prime}}$.